Which of the following is different from the others? (a) 20 + (–25) (b) (– 37) – (– 32) (c) (– 5) × (–1) (d) ( 45 ) ÷ (– 9)

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

Option (c) is correct here
As all the remaining options give a value of -5
20 + (–25) = (–37)–(–32) = ( 45 ) ÷ (–9) = -5
But, (–5) × (–1) = 5



Thanks
Shraddha Singh
2 years ago
Dear Student,
 
20+(-25) = 20-25 = -5
 
(-37)-(-32) = -37+32 = -5
 
(-5)×(-1) = 5
 
45 ÷ (-9) = 45×-1/9 = -5
 
Therefore, in the given four questions (-5)×(-1) is different.
 
Hope it helps!
Thank you